OpenStudy (anonymous):

x-1/9= 2 solve for x

OpenStudy (anonymous):

x-1/9=2 multiply both sides by 9 x-1=18 add 1 to both sides x=19 :D

OpenStudy (mayankdevnani):

x-1/9=2 by cross multiply,